OpenStudy (anonymous):

I need to find the volume of this prism. Can someone please explain how to find it?

OpenStudy (anonymous):

Volume of prism is usually V = LWH

OpenStudy (anonymous):

Depends on what kind of prism tho.

OpenStudy (anonymous):

Volume of triangular prism is V = B * H B is the triangle in this case and H is the side 6.

OpenStudy (anonymous):

You have to find the area of the triangle first though.

OpenStudy (anonymous):

Safe to assume that this is an equilateral triangle so then the triangle ratio is going to 30-60-90.

OpenStudy (anonymous):

That is how you will find the height of the triangle.

OpenStudy (anonymous):

Alright tell me when you find the answer.

OpenStudy (marissalovescats):

Welcome to OpenStudy @rcschooool ! :)

OpenStudy (anonymous):

Thank you I will. I am finding the area of the triangle now

OpenStudy (anonymous):

After you've found the area multiply that by the height (which is 6) and then there is your volume.

OpenStudy (anonymous):

To find the area, I dropped a perpendicular line down to make a right triangle. Then I divided the hypotenuse by 2 to get the short side (which was 3/2) and then used the pythagorean theorem to find the longer side. I got 3/2^2 + b^2 = 3^2 and got 5/2 as the missing side. Is that correct?

OpenStudy (anonymous):

I know that the answer is V= \[something \times \sqrt{3} / 2\] but I do not know what that number is before the sqrt

OpenStudy (anonymous):

@eatacat

OpenStudy (anonymous):

Replying.

OpenStudy (anonymous):



OpenStudy (anonymous):

(2(1/2)1.5*2.59) = 3.885 Area of triangle is 3.885. Volume of prism = B * H = 3.885 * 6 = 23.31

OpenStudy (anonymous):

You already know the hypotenuse since the picture you gave me told me that it was an equilateral triangle. So to find the height you just use the pythagorean formula.

OpenStudy (anonymous):

There is another way to find the height also which is the 30-60-90 I was talking about. If you know your special triangle ratios then you find height. 

OpenStudy (anonymous):

Alright good night.